Homes may be heated by pumping hot water through radiators. What mass of  water will provide the same amount of heat when cooled from 95.0 to 35.0 °C, as  the heat provided when 100 g of steam is cooled from 110 °C to 100 °C.  




1
Expert's answer
2021-12-20T05:00:02-0500

q = mcT. The amount of heat transferred from steam is:

qsteam = 100 g × 1.864 J/(g °C) × (110 – 100) °C = 1864 J

qwater = mwater × 4.184 J/(g °C) × (95 – 35) °C = mwater × 251.04 J/g = qsteam = 1864 J

"m_{water} =\frac{1864J}{251.04J/g} =7.43g"